If the sides of a sqaure are lengthened by 5cm. and the area becomes 81cm^2. What was the length of the original square?
Answer by mananth(16946)

let the length of side be x
x=5 will be the new side
(x+5)^2= 81
take square root of both sides
x+5 =9
x=9-5
x=4 cms.